Could you write this out so I can understand it please.
A Coast Guart patrol boat travels 4 hr on a trip down streem with a 6-mph current. The return trip against the same current takes 5 hr. Find the speed of the boat in still water.

Answer by nerdybill(7384)

A Coast Guart patrol boat travels 4 hr on a trip down streem with a 6-mph current. The return trip against the same current takes 5 hr. Find the speed of the boat in still water.
.
You will be applying the "distance formula":
d = rt
where
d is distance
r is rate or speed
t is time
.
Let x = speed of the boat
then
"distance traveled downstream" = "distance traveled upstream)
4(x+6) = 5(x-6)
4x + 24 = 5x - 30
24 = x - 30
54 mph = x